# A Resource Model for Neural Scaling Laws

### v1

@@ 2 Resource model for single task
$L=\alpha l+\lambda_{1}\theta_{1}+\lambda_{2}\theta_{2},$
where $\alpha$ is a hyperparameter emphasizing the relative importance of task loss compared to regularization. We sweep $\alpha\in[1,512]$. $\theta_{1}\equiv\|\text{\boldmath$\theta$}\|_{1}$ and $\theta_{2}\equiv\|\text{\boldmath$\theta$}\|_{2}$ are L1 and L2 norms of weights, respectively.We fix $\lambda_{1}=0.0001$ and $\lambda_{2}=0.0005$. Such dual regularization approach aims to reduce training variability and foster a clear module structure. The Adam optimizer is used to train the neural network with total epochs set to 100000. The learning rate is initially started at 0.01 and reduced by a factor of 0.3 every 20000 epochs. The batch size is set to 500.
@@ 3 Neuron redundancy in Composite tasks
Resource model hypothesis 2 (Homogeneous growth of resource allocation): If a neural network with $N$ neurons allocates resources $N_{i}$ to subtask $i$, then if the network size is scaled up (along width) such that $N\to aN$, all the resources are scaled up by the same factor $a$, i.e., $N_{i}\to aN_{i}$.
@@ 3.1 Two tasks in parallel
$l=\beta\langle(\hat{y_{1}}-y_{1})^{2}\rangle+(1-\beta)\langle(\hat{y_{2}}-y_{2})^{2}\rangle,$
@@
$L=\alpha l+\lambda_{1}\theta_{1}+\theta_{2}\ell_{2}.$
@@
[…]
@@
$l=\langle(\hat{y}-\sqrt{(x_{1}-x_{2})^{2}+(x_{3}-x_{4})^{2}})^{2}\rangle,$
@@
$L=\alpha l+\lambda_{1}\theta_{1}+\lambda_{2}\theta_{2}.$
@@
[…]
@@ 3.3 Conjecture for general composite tasks
Resource model hypothesis 3 (Linear additivity of losses): The loss of a composite task can be decomposed as a linear summation of the losses of its subtasks.
@@ B.2 Matching Ensembling mechansim with $x^{2}$ regression experiment
We revealed that the ensembling mechanism, suggested in  Liu & Tegmark (2023), does not convincingly account for loss scaling observed in a toy experiment for single $x^{2}$ regression task. In the experiment, ensembling is expected to occur between outputs from several $x^{2}$ regressors, pairs of symmetric neurons that have the same bias and the opposite weight in the first layer  Liu & Tegmark (2023). We identified multiple $x^{2}$ regressors through matching pair of neurons by minimizing relative distances when projected onto their counterpart as in Figure 6(left). We ploted the error functions(center) across the input distribution and their correlations(right). Correlation values are ±$\pm 1$ in most cases, indicating these error function are highly correlated(anticorrelated). Therefore, we concluded that ensembling does not occur for a set of $x^{2}$ regressors in our single ${}^{2}$ regression toy experiment.

### v2

@@ 2 Resource model for single task
$L=\alpha l+\lambda_{1}\sum\|\theta_{w}\|+\lambda_{2}\sum{\|\theta_{w}}\|^{2}$
where $\alpha$ is a hyperparameter emphasizing the relative importance of task loss compared to regularization. $\sum\|\theta_{w}\|$ and $\sum{\|\theta_{w}}\|^{2}$ are L1 and L2 norms of weights, respectively. We sweep $\alpha\in[1,512]$. We fix $\lambda_{1}=0.0001$ and $\lambda_{2}=0.0005$. Such dual regularization approach aims to reduce training variability and foster a clear module structure. The Adam optimizer is used to train the neural network with total epochs set to 100000. The learning rate is initially started at 0.01 and reduced by a factor of 0.3 every 20000 epochs. The batch size is set to 500.
@@ 3 Neuron redundancy in Composite tasks
Resource model hypothesis 2 (Homogeneous growth of neuron allocation): If a neural network with $N$ neurons allocates resources $N_{i}$ to subtask $i$, then if the network size is scaled up (along width) such that $N\to aN$, all the resources are scaled up by the same factor $a$, i.e., $N_{i}\to aN_{i}$.
@@ 3.1 Two tasks in parallel
$l=\beta\langle(\hat{y_{1}}-y_{1})^{2}\rangle+(1-\beta)\langle(\hat{y_{2}}-y_{2})^{2}\rangle$
@@
$L=\alpha l+\lambda_{1}\sum\|\theta_{w}\|+\lambda_{2}\sum{\|\theta_{w}}\|^{2}$
@@
Consider a set of target functions $f_{i}:x\rightarrow f_{i}(x)$ for $i=1,2,\ldots,Q$, and corresponding number of allocated neurons $N_{i}$. Consider a composite task, which is a linear and parallel combination of these regression tasks i.e composite task loss is $l=\sum_{i\in Q}\alpha_{i}{\rm MSE}_{i}$. Assuming a fixed relative ratio of resources for each subtask (the homogeneous growth hypothesis), the composite loss scales as $\sim N_{\text{total}}^{-1}$, where $N_{\text{total}}=\sum_{i}N_{i}$ is the total number of neurons allocated to the composite task.
@@
$l=\langle(\hat{y}-\sqrt{(x_{1}-x_{2})^{2}+(x_{3}-x_{4})^{2}})^{2}\rangle$
@@
$L=\alpha l+\lambda_{1}\sum\|\theta_{w}\|+\lambda_{2}\sum{\|\theta_{w}}\|^{2}$
@@
: Consider a set of target functions $f_{i}:x\rightarrow f_{i}(x)\ (i=1,2,\ldots,Q)$. Consider a composite task, which is a serial combination of such regression tasks i.e. $l=MSE(\hat{y},(f_{1}\circ f_{2}\circ f_{3}\circ\ldots\circ f_{Q})(x))$. Assuming a fixed relative ratio of resources for each subtask (homogeneous growth of resource allocation) and a linear seperability of the composite task loss function, then the total loss scales $\sim N_{\text{total}}^{-1}$, where $N_{\text{total}}=\sum_{i}N_{i}$ is the total number of neurons allocated to the composite task.
@@ 3.3 Conjecture for general composite tasks
Resource model hypothesis 3 (Linear additivity of subtasks losses): The loss of a composite task can be decomposed as a linear summation of the losses of its subtasks.
@@ B.2 Matching Ensembling mechansim with $x^{2}$ regression experiment
We revealed that the ensembling mechanism, suggested in  Liu & Tegmark (2023), does not convincingly account for loss scaling observed in a toy experiment for single $x^{2}$ regression task. In the experiment, ensembling is expected to occur between outputs from several $x^{2}$ regressors, pairs of symmetric neurons that have the same bias and the opposite weight in the first layer  Liu & Tegmark (2023). We identified multiple $x^{2}$ regressors through matching pair of neurons by minimizing relative distances when projected onto their counterpart as in Figure 6(left). We ploted the error functions(center) across the input distribution and their correlations(right). Correlation values are ±$\pm 1$ in most cases, indicating these error function are highly correlated(anticorrelated). Therefore, we concluded that ensembling does not occur for a set of $x^{2}$ regressors in our single 2 regression toy experiment.